a body is dropped from a certain height H if the ratio oh the distances travelled by it in (n-3) seconds and (n-3)th second is 4:3 find H (g=10)

Grade:9

1 Answers

Sanjeev Malik IIT BHU
12 years ago

let n-3 = t. as dropped initial speed =0.

dis in t sec , s= gt2/2     ;  dis in t+1 sec = g(t+1)2/2

so dis in tth sec = g(t+1)2/2- gt2/2 =  gt+ g/2

4/3= (gt2/2)/(gt +g/2) =t2/(2t+1)

find the value of t and then n.

then height H = gn2/2